OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, five divided by three, negative five divided by eighteen, ...

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf \textit{sum of an infinite geometric serie}=\Sigma_{i=0}^\infty\quad a_1\cdot r^i=\cfrac{a_1}{1-r}\) so... .we know the 1st term term... is there.. just 60 to find "r" in a geometric sequence... just divide the "next term" by the "current term" and that will spit out the multiplier "r"

OpenStudy (jdoe0001):

convergent... means it ends somewhere at some number divergent... means.. it doesn't have a limit, it just keep on going

OpenStudy (jdoe0001):

so.. get "r" and plug it in at \(\large \bf \cfrac{a_1}{1-r}\) if it gives a value, that's it's limit, it ends there, thus is convergent

OpenStudy (anonymous):

r=-6?

OpenStudy (jdoe0001):

well.. not quite. close

OpenStudy (jdoe0001):

\(\bf \cfrac{-\cancel{ 10 }}{\cancel{ 60 }}=?\)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

is it positive 6?

OpenStudy (anonymous):

-1/6?

OpenStudy (jdoe0001):

hehe... yes.. -1/6 so there :)... just plug that in

OpenStudy (jdoe0001):

\(\bf \cfrac{60}{1-\left(-\frac{1}{6}\right)}\)

OpenStudy (anonymous):

360/7

OpenStudy (jdoe0001):

yeap.... so it ends somewhere.. thus is convergent

OpenStudy (anonymous):

wheres the limit?

OpenStudy (jdoe0001):

360/7 <--

OpenStudy (anonymous):

hmm, that's not of my answer choices :/

OpenStudy (jdoe0001):

hmm 51.43 I'd think is the decimal of that

OpenStudy (anonymous):

yes, but it is still not of my answer choices.

OpenStudy (jdoe0001):

hmmm what are the choices given? I don't see anything wrong there

OpenStudy (anonymous):

Diverges Converges; 11100 Converges; 72 Converges; 0

OpenStudy (anonymous):

i think r was suppose to be positive, 1/6, so we get the limit 72. we are multiplying 1/6 everytime.

OpenStudy (jdoe0001):

hmmm ....

OpenStudy (jdoe0001):

so....... can't be.... is a negative 2nd and rth term... so is a negative "r"

OpenStudy (jdoe0001):

2nd and 4th term I meant.... but.... hmmm nope.. looks to me ok with 360/7

OpenStudy (jdoe0001):

maybe the negatives aren't meant to be there... or someone forgot to add the minus to the 1/6 when getting the choices done

OpenStudy (anonymous):

i will ask my teacher about it.

OpenStudy (jdoe0001):

it happens =)

OpenStudy (anonymous):

everyone makes mistakes C;